OpenStudy (anonymous):

Please help with this! List the coordinates of the vertex & list two other points on the parabola: x=y^2-2

OpenStudy (anonymous):

Put the equation in vertex form first

OpenStudy (anonymous):

How do i do that?

OpenStudy (anonymous):

Get variables on one side and constant on the other

OpenStudy (anonymous):

then complete the square

OpenStudy (anonymous):

I dont understand any of this. what square?

OpenStudy (anonymous):

would the vertex be (0,-2) ?

OpenStudy (anonymous):

ok. x=y^2-2 so you can pick a point for x and solve for y y=\[\sqrt{x+2}\] this will give you a few points on the parabola

OpenStudy (anonymous):

so if x=1, then y=root 3, etc.

OpenStudy (anonymous):

would one of the points be (-1.5,0) ?

OpenStudy (anonymous):

The x value of the vertex is -b/2a from the quadratic formula then substitute to find the y value f(-b/2a)

OpenStudy (anonymous):

if you put in -1.5 you get sqroot of .5 so no

OpenStudy (anonymous):

well would the vertex be (-2,0) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, (: now can you help me find two points?

OpenStudy (anonymous):

yes. Pick two numbers for x; plug them in the equation and solve for y. y=root(x+2)